OpenStudy (anonymous):

The time (in years) after reaching age 60 that it takes an individual to retire is approximately exponentially distributed with a mean of about 6 years. Suppose we randomly pick one retired individual. We are interested in the time after age 60 to retirement. Find the probability that the person retired after age 68.

OpenStudy (anonymous):

Jesus. I once knew how to do that kind of stuff. Tell me the name of that thing your studying so i can remember the formula. Is it poisson distribution, ANOVA, Variance, ... ?

OpenStudy (anonymous):

Um, Exponential Distribution? At least, that's what chapter it was in.

OpenStudy (anonymous):

Oh, this is also continuous random variables instead of discrete random variables.

OpenStudy (anonymous):

Ok so now i have a problem... I dont know if i only use values from 60+ or if i have to use the real number as the text says: We are interested in the time after age 60 to retirement. The 2 option: as the mean = 6 and \[mean = 1 / \lambda \] then \[\lambda\]= 1/6 P(X) = \[\lambda*e ^{-\lambda*X}\] P(8) = (1/6)*e^(-(1/6)*8) = 4.4% OR mean = 66 P(68) = 0.54% But i think the correct answer is the first one, 4.4%